### v4

# Imprecise Markov Semigroups and their Ergodicity

Abstract.
We introduce the concept of an imprecise Markov semigroup $\mathbf{Q}$. It is a tool that allows us to represent ambiguity around both the initial and the transition probabilities of a continuous-time Markov process via a compact collection of Markov semigroups, each associated with a (possibly different) Markov process. We use techniques from topology, geometry, and probability to study the ergodic behavior of $\mathbf{Q}$. We show that, under some conditions that also involve the geometry of the state space, eventually the ambiguity fades. We call this property ergodicity of the imprecise Markov semigroup, and we relate it to the classical notion of ergodicity. We prove ergodicity both when the state space is Euclidean or a Riemannian manifold, and when it is an arbitrary measurable space. The importance of our findings for the fields of artificial intelligence and computer vision is also discussed, in particular in the study of how the probability of an output evolves over time as we perturb the input of a convolutional autoencoder.
Key words and phrases:
Imprecise Markov processes, Markov Diffusion Operators, Bakry-Émery curvature, Ergodic Processes, Autoencoders
1991 Mathematics Subject Classification:
60J60, 58J65, 62A01
1. Introduction
When dealing with a stochastic phenomenon whose future evolution is independent of its history, it is customary to model its behavior via a stochastic process enjoying the Markov property. Loosely, this means that the law of the realization $X_{T}$ of the process $(X_{t})_{t\geq 0}$ at time $T$, given what happened up to time $s<T$, is equivalent to the law of $X_{T}$, given only $X_{s}$. We can of course express this property in a slightly more general way, by saying that the expectation of a generic bounded measurable real-valued function $\tilde{f}$ of $X_{T}$, given what happened in the past up to time $s<T$, is equal to the expectation of $\tilde{f}(X_{T})$, given only $X_{s}$. [^1] Throughout the paper, we do not use the pronoun “we” as a form of pluralis majestatis, but rather to comply to the convention of scientific papers of writing in plural form. Notice that in the present paper we perpetrate a terminology abuse, and refer to real-valued functions on the state space $E$ as “functionals”. In addition, we only consider continuous-time processes $(X_{t})_{t\geq 0}$.
Two elements play a key role in a process enjoying the Markov property. The first is the law of the initial state $X_{0}$, and the second is the transition probability from the initial state $X_{0}$ to any state $X_{t}$ of the process, i.e. the probability that, given that the process begins at $X_{0}$, it “reaches a certain value” at time $t$. In many applications, relaxing the hypothesis that one or both of them are precise is natural, and leads to less precise but more robust results.
Imprecise probability (Augustin et al., 2014; Troffaes and de Cooman, 2014; Walley, 1991) is a field whose techniques naturally lend themselves to be used for this purpose. In fact, imprecise Markov processes are a flourishing area of the imprecise probability literature (Crossman and Škulj, 2010; De Bock et al., 2021; de Cooman et al., 2009; Delgado et al., 2011; Jaeger et al., 2020; Krak et al., 2017, 2019; T’Joens and De Bock, 2021; Trevizan et al., 2007, 2008; Troffaes et al., 2019), [^2] Much of the existing literature focuses on discrete-time processes. with numerous applications e.g. in engineering (Aslett et al., 2022, Chapter 4), including aerospace engineering (Vasile, 2021, Chapter 5), and economics (Cerreia-Vioglio et al., 2023; Denk et al., 2018, 2020). As the name suggests, imprecise Markov processes allow to take into account imprecision in the initial law and the transition probabilities of a Markov process.
In this paper, we introduce a new concept that we call an Imprecise Markov Semigroup. It is a compact collection of (precise) Markov semigroups, each associated with a Markov process. [^3] Here, by “precise” we refer to the classical notion of a Markov semigroup. Modeling the aforementioned types of imprecision via such a collection, instead of using lower previsions (Troffaes and de Cooman, 2014) or Choquet integrals (Choquet, 1954), allows us to develop new geometric and probabilistic techniques based on Bakry et al. (2014) that make the study of the limiting behavior of imprecise Markov processes (captured by the limit behavior of the imprecise Markov semigroup) easier. [^4] Since we consider a collection of precise Markov semigroups, one could say ours is a “precise approach” to imprecision. In particular, we discover a relationship between such a behavior and the curvature of the state space – that is, the space $E$ that the elements of the process $(X_{t})_{t\geq 0}$ take values on. This also relates the present work to the study of the geometry of uncertainty (Cuzzolin, 2021).
Our goal is to study what we call the $\tilde{f}-$ergodic property of an imprecise Markov semigroup. Roughly, we want to see whether a form of the classical ergodic mantra “in the limit, time average (of the form $\lim_{t\rightarrow\infty}\mathbb{E}[\tilde{f}(X_{t})\mid X_{0}]$) equals space average (of the form $\mathbb{E}[\tilde{f}(X_{0})]$)” can be derived (Cerreia-Vioglio et al., 2015; Caprio and Mukherjee, 2023), for some bounded functional $\tilde{f}$ of interest. Slightly more precisely, we inspect whether the limits as $t\rightarrow\infty$ of the maximum and the minimum values for the expectation $\mathbb{E}[\tilde{f}(X_{t})\mid X_{0}]$ – captured by the maximum and minimum elements of the imprecise Markov semigroup – are equal to the expectation $\mathbb{E}[\tilde{f}(X_{0})]$, taken with respect to some measures of interest. The answer, which – as we shall see – follows from Corollary 8.1 (where the state space $E$ is a Euclidean space or a Riemannian manifold) and Corollary 10.1 (where $E$ is an arbitrary measurable space), is positive, under some conditions (including some geometric ones, as we briefly mentioned earlier).
In addition, we discover that – under a further condition interpretable as the initial distribution of the process, i.e. the distribution of $X_{0}$, being known and invariant – in the limit the imprecision vanishes. We find that the limits for both the maximum and the minimum of $\mathbb{E}[\tilde{f}(X_{t})\mid X_{0}]$ coincide, and are equal to $\mathbb{E}[\tilde{f}(X_{0})]$ (taken with respect to the initial distribution of the process). The similarities and differences between our results and the existing ones in the imprecise Markov literature, and in imprecise ergodic theory (Caprio and Gong, 2023; Caprio and Mukherjee, 2023; Cerreia-Vioglio et al., 2015), are discussed in Sections 2 and 3.
Besides being interesting in their own right (and because they provide new techniques to the fields of imprecise Markov processes), we believe that our findings will prove useful in future research in machine learning, artificial intelligence, and computer vision. Indeed, in recent works on convolutional autoencoders, images (or, more in general, inputs from an input space $\mathcal{X}$) are “projected” onto a low-dimensional manifold $E$ by an encoding function $\phi_{\text{enc}}:\mathcal{X}\rightarrow E$. Different “portions” of the surface $\mathcal{S}(E)$ of such a manifold capture different features of the input image (Yu et al., 2023). For instance, a portion of the surface may correspond to the feature “dog”, so that dog images are projected there. This means that we can consider a partition $\mathscr{E}=\{E_{j}\}_{j=1}^{J}$ of $\mathcal{S}(E)$ whose elements correspond to “feature portions”. For example, $J$ could be equal to the cardinality $|\mathcal{Y}|$ of the output space $\mathcal{Y}$ in the case of classification problems. Then, a decoding function $\phi_{\text{dec}}:E\rightarrow\text{Prob}(\mathcal{Y})$ returns the probabilities of $y_{j}\in\mathcal{Y}$ being the correct output for the input $x\in\mathcal{X}$. Here $\text{Prob}(\mathcal{Y})$ denotes the space of countably additive probabilities on $\mathcal{Y}$. For example, in a classification case, $(\phi_{\text{dec}}\circ\phi_{\text{enc}})(x)=(P(y_{1}),\ldots,P(y_{J}))^{\top}$, where $\mathcal{Y}=\{y_{1},\ldots,y_{J}\}$ is the label space, and $P(y_{j})$ is the probability estimated by the model that $y_{j}$ is the correct label for input $x$, $j\in\{1,\ldots,J\}$. A visual representation is given in Figure 1.
Figure 1. In this figure, we provide a visual representation of the functioning of a convolutional autoencoder in an image classification problem. The input image of a pug is mapped by encoding function $\phi_{\text{enc}}$ to the surface area $E_{\text{dog}}\in\mathscr{E}$ corresponding to the feature “dog”. Such a projection is then mapped by decoding function $\phi_{\text{dec}}$ to the probability space over the labels. In this simple example the model gives a high probability to the first label, i.e. “dog”.
Now, imagine that we want to study a smooth transition between a dog and a cat image. Intuitively, this corresponds to a “walk” from the “dog portion” $E_{\text{dog}}$ of the manifold’s surface $\mathcal{S}(E)$ to the “cat portion” $E_{\text{cat}}$ (Pérez, 1998). A visual example is given in Figure 2. Our results, and especially Corollary 8.1, may be very useful when (i) such a “walk” is not deterministic, but rather modeled via a Markov process; (ii) the user is uncertain about which initial and transition probabilities to choose; and (iii) we are interested in the probability of a particular label during such a walk. For example, if we are interested in how the probability of the first label evolves over time, we could consider function $z\mapsto\tilde{f}(z)=e_{1}^{\top}\phi_{\text{dec}}(z)$, where $e_{1}=(1,0,0,\ldots,0)^{\top}$ is a $J$-dimensional standard basis vector, and $\tilde{f}$ is an example of a bounded functional on $E$, as discussed earlier in this section.
Figure 2. In this figure, we represent the smooth transition between a dog and a cat pictures. Moving a cursor between the pictures causes a walk from the “dog portion” $E_{\text{dog}}$ of the manifold’s surface $\mathcal{S}(E)$ to the “cat portion” $E_{\text{cat}}$, depicted as a purple smooth curve. The results we present in this paper are useful to determine the behavior of a similar walk that is not deterministic, when it is not possible to determine unique initial and transition probabilities, and when we are interested e.g. in the evolution of the probability of the first label, $y_{1}=\text{``dog''}$, captured by the functional $z\mapsto\tilde{f}(z)=e_{1}^{\top}\phi_{\text{dec}}(z)$.
Our paper is structured as follows. Section 2 introduces the concept of an Imprecise Markov Semigroup. Section 3 derives our results when state space $E$ is a Euclidean space or a Riemannian manifold. Section 4 generalizes the findings in Section 3 to the case of $E$ being an arbitrary measurable space. Section 5 concludes our work. We prove Theorem 10 in A, and we provide a primer on Markov semigroups and diffusion Markov triples in B.
2. Imprecise Markov Semigroups
In this section, we introduce the concept of an Imprecise Markov Semigroup. Its definition hinges on the notions of Markov processes, Markov semigroups, carré du champ and iterated carré du champ operators, and Bakry-Émery curvature. We refer the unfamiliar reader to B.1.
The following is introduced in Bakry et al. (2014, Chapter 1.1). It is the minimal structure we require throughout the paper for the space where our stochastic processes of interest live.
Definition 1 (Good Measurable Space (Bakry et al., 2014)).
A good measurable space is a measurable space $(E,\mathcal{F})$ for which the measure decomposition theorem applies, and for which there is a countable family of sets which generates the $\sigma$-algebra $\mathcal{F}$.
In this work, we are interested in Markov semigroups $\mathbf{P}=(P_{t})_{t\geq 0}$ associated with (time homogeneous) Markov processes $(X_{t})_{t\geq 0}$. For such a Markov semigroup, for every $t\geq 0$, operator $P_{t}$ is such that
$P_{t}f(x)=\mathbb{E}\left[f(X_{t})\mid X_{0}=x\right]=\int_{E}f(y)p_{t}(x,\text{d}y),\quad f\in B(E)\text{, }x\in E,$
where $B(E)$ is the space of bounded measurable functionals on $E$, and $p_{t}$ is the probability kernel describing the evolution of the Markov process. In words, operator $P_{t}$ applied on $f\in B(E)$ gives us the expected value of $f(X_{t})$, “weighted by” the transition probability (from the starting value $X_{0}=x$ to $X_{t}$) of our Markov process. We focus on this type of Markov semigroups because studying their (ergodic) behavior inform us about the (ergodic) behavior of the associated Markov processes.
Throughout the paper, we will assume that the semigroups we consider always have an invariant probability measure $\mu$, so that the elements of $\mathbf{P}=(P_{t})_{t\geq 0}$ are bounded operators in $\mathbb{L}^{p}(\mu)$, $p\in[1,\infty)$. In fact, most semigroups of interest have an invariant measure (Bakry et al., 2014, Section 1.2.1). A way to derive $\mu$ as a weak limit of a “reasonable” initial probability measure $\mu_{0}$ is inspected in Bakry et al. (2014, Page 10). For a discussion on this matter, we refer the interested reader to Remark 2 in B.1.
We now introduce a space that will play a pivotal role in the definition of an Imprecise Markov Semigroup. Recall that, for all $t\geq 0$, the $t$-th element $P_{t}$ of Markov semigroup $\mathbf{P}=(P_{t})_{t\geq 0}$ is a linear operator that sends bounded measurable functionals on $E$ to bounded measurable functionals on $E$. In formulas, we can write this as $P_{t}:B(E)\rightarrow B(E)$, or equivalently, $P_{t}\in B(E)^{B(E)}$. In turn, this implies that a Markov semigroup $\mathbf{P}$ can be seen as a function mapping bounded measurable functionals $f$ on $E$ to nets of bounded measurable functionals $\mathbf{P}f=(P_{t}f)_{t\geq 0}$ on $E$. In formulas, $\mathbf{P}:B(E)\rightarrow{B(E)}^{\mathbb{R}_{+}}$, or equivalently, $\mathbf{P}\in({B(E)}^{\mathbb{R}_{+}})^{B(E)}$.
Consider now a collection $\mathbf{Q}$ of Markov semigroups (MSGs) associated with Markov processes. That is, let $\mathbf{Q}\subset({B(E)}^{\mathbb{R}_{+}})^{B(E)}$ be such that
[…]
The notion of Imprecise Markov Semigroup hinges on the possibility of comparing the elements of $\mathbf{Q}$, when they are evaluated at some functional $\tilde{f}\in B(E)$. First, suppose we are able to define a total order $\preceq^{\text{tot}}_{\tilde{f}}$ so that
$\mathbf{P}\preceq^{\text{tot}}_{\tilde{f}}\mathbf{P^{\prime}}\iff P_{t}\tilde{f}\leq P^{\prime}_{t}\tilde{f}\text{, }\forall t\geq 0.$
Being $\preceq^{\text{tot}}_{\tilde{f}}$ a total order, this means that, for all $\mathbf{P},\mathbf{P^{\prime}}\in\mathbf{Q}$, either $\mathbf{P}\preceq^{\text{tot}}_{\tilde{f}}\mathbf{P^{\prime}}$ or $\mathbf{P^{\prime}}\preceq^{\text{tot}}_{\tilde{f}}\mathbf{P}$. In addition, the notions of maximal and greatest elements of $(\mathbf{Q},\preceq^{\text{tot}}_{\tilde{f}})$ coincide, and correspond to the classical notion of maximum (with respect to the total order $\preceq^{\text{tot}}_{\tilde{f}}$). Similarly, minimal and least element notions coincide, and correspond to the classical notion of minimum.
Now, let $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$ be the order topology. That is (provided $|\mathbf{Q}|\geq 2$), $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$ is the topology whose base is given by the open rays $\{\mathbf{P}:\mathbf{P}_{1}\prec^{\text{tot}}_{\tilde{f}}\mathbf{P}\}$ and $\{\mathbf{P}:\mathbf{P}\prec^{\text{tot}}_{\tilde{f}}\mathbf{P}_{2}\}$, for all $\mathbf{P}_{1}$, $\mathbf{P}_{2}$ in $\mathbf{Q}$, and by the open intervals $(\mathbf{P}_{1},\mathbf{P}_{2})=\{\mathbf{P}:\mathbf{P}_{1}\prec^{\text{tot}}_{\tilde{f}}\mathbf{P}\prec^{\text{tot}}_{\tilde{f}}\mathbf{P}_{2}\}$. This means that the $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$-open sets in $\mathbf{Q}$ are the sets that are a union of (possibly infinitely many) such open intervals and rays.
Assume that $\mathbf{Q}$ is $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$-compact. This implies that $\mathbf{Q}$ contains the minimum and the maximum (with respect to the total order $\preceq^{\text{tot}}_{\tilde{f}}$), which we denote by $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$, respectively. In turn, since $\underline{\mathbf{P}},\overline{\mathbf{P}}\in\mathbf{Q}$, they both are (precise) Markov semigroups associated with Markov processes by (1).
Suppose now that – contrary to before – we are only able to define a partial order $\preceq^{\text{part}}_{\tilde{f}}$ on $\mathbf{Q}$ as the previous one. More formally, we have that
$\mathbf{P}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime}}\iff P_{t}\tilde{f}\leq P^{\prime}_{t}\tilde{f}\text{, }\forall t\geq 0,$
but there might exist $\mathbf{P},\mathbf{P^{\prime}}\in\mathbf{Q}$ such that $\mathbf{P}\not\preceq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime}}$ and $\mathbf{P}\not\succeq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime}}$. Let us now introduce three set theoretic concepts (Wolk, 1958), that will allow us to properly define an Imprecise Markov Semigroup even when only a partial order can be defined.
Definition 2 (Up/Down-Directed Subsets of $\mathbf{Q}$).
A subset $\mathbf{S}$ of $\mathbf{Q}$ is up-directed if for all $\mathbf{P},\mathbf{P^{\prime}}\in\mathbf{S}$, there exists $\mathbf{P^{\prime\prime}}\in\mathbf{S}$ such that $\mathbf{P}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime\prime}}$ and $\mathbf{P^{\prime}}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime\prime}}$. We say that $\mathbf{S}$ is down-directed if the opposite holds. That is, if for all $\mathbf{P},\mathbf{P^{\prime}}\in\mathbf{S}$, there exists $\mathbf{P^{\prime\prime}}\in\mathbf{S}$ such that $\mathbf{P^{\prime\prime}}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P}$ and $\mathbf{P^{\prime\prime}}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime}}$.
Definition 3 (Dedekind-Closed Subsets of $\mathbf{Q}$).
A subset $\mathbf{K}$ of $\mathbf{Q}$ is Dedekind-closed if whenever $\mathbf{S}$ is an up-directed subset of $\mathbf{K}$, and $\mathbf{P}$ is the $\preceq^{\text{part}}_{\tilde{f}}$-least upper bound of $\mathbf{S}$ (or $\mathbf{S}$ is a down-directed subset of $\mathbf{K}$, and $\mathbf{P}$ is the $\preceq^{\text{part}}_{\tilde{f}}$-greatest lower bound of $\mathbf{S}$), then $\mathbf{P}\in\mathbf{K}$.
Definition 4 (Incomparability, Diversity, and $\preceq^{\text{part}}_{\tilde{f}}$-Width of $\mathbf{Q}$).
[…]
Definitions 2, 3, and 4 are needed to introduce the concept of $\preceq^{\text{part}}_{\tilde{f}}$-compatible topology. Requiring that $\mathbf{Q}$ is compact in such a topology (together with another assumption) will ensure us that it possesses least and greatest elements (with respect to the partial order $\preceq^{\text{part}}_{\tilde{f}}$).
Definition 5 ($\preceq^{\text{part}}_{\tilde{f}}$-Compatible Topology (Wolk, 1958)).
A topology $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$ on $\mathbf{Q}$ is $\preceq^{\text{part}}_{\tilde{f}}$-compatible if
  • every $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-closed set is also Dedekind-closed
  • every set of the form $\{\mathbf{P}\in\mathbf{Q}:\mathbf{P^{\prime}}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P^{\prime\prime}}\}$ is $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-closed.
Lemma 6 (Topological Properties of $\mathbf{Q}$).
If $\mathbf{Q}$ has finite $\preceq^{\text{part}}_{\tilde{f}}$-width, then $\preceq^{\text{part}}_{\tilde{f}}$-compatible topology $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$ is unique. If in addition $\mathbf{Q}$ is $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compact, then it possesses least and greatest elements (with respect to the partial order $\preceq^{\text{part}}_{\tilde{f}}$). We denote them by $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$, respectively.
Proof.
The first part of the lemma is a consequence of Wolk (1958, Theorem, page 528), which also shows that $(\mathbf{Q},\tau_{\preceq^{\text{part}}_{\tilde{f}}})$ is Hausdorff. The second part is an immediate consequence of requiring that $\mathbf{Q}$ is $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compact. Indeed, from Definition 5, this entails that $\mathbf{Q}$ possesses least and greatest elements. ∎
Let us add a discussion on why the conditions in Lemma 6 ensure us that the greatest and least elements for $\mathbf{Q}$ exist. Compactness with respect to the topology $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$ means that the limit points of the chains [^5] Recall that a chain in $\mathbf{Q}$ is a subset of $\mathbf{Q}$ that is totally ordered with respect to $\preceq^{\text{part}}_{\tilde{f}}$}. in the partially ordered set (poset) $\mathbf{Q}$ exist, and belong to the poset $\mathbf{Q}$ itself. In other words, the $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compactness condition ensures that the poset $\mathbf{Q}$ does not “escape to infinity” in a topological sense.
Finite $\preceq^{\text{part}}_{\tilde{f}}$-width restricts the degree of “spread” in the poset $\mathbf{Q}$, preventing it from having arbitrarily large levels of incomparability. In other words, it ensures us that the “pathological elements” of $\mathbf{Q}$ – i.e. the incomparable ones – are controllably few. This helps us closing the gap with respect to the total order $\preceq^{\text{tot}}_{\tilde{f}}$ case, for which no incomparable elements exist.
Let us add a simple example of a collection $\mathbf{Q}$ that satisfies (1) and has finite $\preceq^{\text{part}}_{\tilde{f}}$-width. Consider a finite collection $\mathbf{Q}=\{\mathbf{P}_{i}\}_{i=1}^{6}$ of six Markov semigroups $\mathbf{P}_{i}$ associated with as many Markov processes. Fix a functional $\tilde{f}\in B(E)$, and suppose that the partial order $\preceq^{\text{part}}_{\tilde{f}}$ can be defined, so that we can draw the following Hasse diagram
$\mathbf{P}_{6}$$\mathbf{P}_{2}$$\mathbf{P}_{3}$$\mathbf{P}_{4}$$\mathbf{P}_{5}$$\mathbf{P}_{1}$
Notice that, from top to bottom, a line segment represents the concept “is preferred to according to the partial order $\preceq^{\text{part}}_{\tilde{f}}$”, and if no such a segment connects two elements, this means that they are not comparable. So, for instance, $\mathbf{P}_{2}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P}_{6}$ and $\mathbf{P}_{1}\preceq^{\text{part}}_{\tilde{f}}\mathbf{P}_{2}$, but $\mathbf{P}_{2}$ and $\mathbf{P}_{3}$ are incomparable. Then, the $\preceq^{\text{part}}_{\tilde{f}}$-width of $\mathbf{Q}$ is $4$, since the diverse set having the largest cardinality is $\mathbf{S}=\{\mathbf{P}_{2},\mathbf{P}_{3},\mathbf{P}_{4},\mathbf{P}_{5}\}$. In addition, we see that $\mathbf{Q}$ has a least element $\mathbf{P}_{1}$ and a greatest element $\mathbf{P}_{6}$; it is also trivially $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compact since it is finite.
Together, $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compactness and finite $\preceq^{\text{part}}_{\tilde{f}}$}-width ensure the existence of the greatest and least elements $\overline{\mathbf{P}}$ and $\underline{\mathbf{P}}$, respectively, in the poset $\mathbf{Q}$. We are now ready for the full definition of an Imprecise Markov Semigroup.
Definition 7 (Imprecise Markov Semigroup).
Let $\mathbf{Q}$ be a subset of the space $({B(E)}^{\mathbb{R}_{+}})^{B(E)}$ that satisfies (1). Fix any $\tilde{f}\in B(E)$. If we can define the total order $\preceq^{\text{tot}}_{\tilde{f}}$ on $\mathbf{Q}$, then we say that $\mathbf{Q}$ is an Imprecise Markov Semigroup if it is $\tau_{\preceq^{\text{tot}}_{\tilde{f}}}$-compact.
If instead we can only specify the partial order $\preceq^{\text{part}}_{\tilde{f}}$, then we say that $\mathbf{Q}$ is an Imprecise Markov Semigroup if it has finite $\preceq^{\text{part}}_{\tilde{f}}$-width, and is $\tau_{\preceq^{\text{part}}_{\tilde{f}}}$-compact.
In both cases, we denote by $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$ its least and greatest elements, respectively.
Let us add two remarks. First, the reader may ask themselves why we went to such great lengths to ensure that $\mathbf{Q}$ possesses the least and greatest elements $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$, respectively. Informally, the reason is that studying their limiting behavior (that is, studying the limits $\lim_{t\rightarrow\infty}\underline{P}_{t}\tilde{f}$ and $\lim_{t\rightarrow\infty}\overline{P}_{t}\tilde{f}$) will inform us on the limiting behavior of all the elements of $\mathbf{Q}$ evaluated at $\tilde{f}$. This intuition is made formal in Corollaries 8.1 and 10.1. In addition, since $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$ are (precise) Markov semigroups, we can use techniques developed e.g. in Bakry et al. (2014) to carry out such a study.
[…]
It is also worth mentioning that we focus our attention on orders (total and partial) associated with a particular function $\tilde{f}$ because, given an order of the type $\mathbf{P}\preceq\mathbf{P^{\prime}}\iff P_{t}f\leq P^{\prime}_{t}f$, for all $t\geq 0$ and all $f\in B(E)$, there exists at least one set $\mathbf{Q}$ that satisfies (1), but which does not possess a least element $\underline{\mathbf{P}}$, according to such an order (Krak et al., 2017, Example 6.2).
As we can see, the choice of $\tilde{f}$ is then crucial, and it should be driven by the application of interest. For example, as we discussed in Section 1, in the case of a convolutional autoencoder, we can pick $\tilde{f}(z)=e_{1}^{\top}\phi_{\text{dec}}(z)$ to study how the probability of the first label evolves over time. In quantitative finance, we could pick $\tilde{f}(z)=\min\{(K-z_{i})_{+},M\},$ $z\in\mathbb{R}_{+}^{d}$, where $(\cdot)_{+}=\max(\cdot,0)$, $K$ is the strike, and $M$ a truncation level. This is a truncated put‑option payoff on the $i$-th asset; the resulting lower and upper semigroups yield worst‑case option prices under model uncertainty. In reliability/survival analysis, a viable choice is $\tilde{f}(z)=\text{Ind}_{\{z\leq\theta\}},$ for a failure threshold $\theta$, where $\text{Ind}_{\{\cdot\}}$ denotes the indicator function. The functional tracks the probability that the degradation index $X_{t}$ remains in the “healthy” region, yielding robust upper and lower survival curves.
[…]
Let us pause here to expand on the last point. Working with Imprecise Markov Semigroups (IMSGs) allows us to capture the ambiguity faced by the agent in assessing both the invariant measure and the transition probability of a Markov process. By studying IMSGs, we take a different perspective to imprecise Markov processes than the ones usually employed in the literature introduced in Section 1. Indeed, existing works model invariant measure ambiguity via an invariant Choquet capacity – see e.g. de Cooman et al. (2009) – and transition probability ambiguity via sets of transition probabilities or sets of transition rate matrices, usually assumed convex to allow maximization and minimization operators to be computed in terms of linear programming optimization problems – see e.g. Krak et al. (2017, 2019). In this paper, instead, we capture both types of ambiguities via the imprecise Markov semigroup $\mathbf{Q}$. From Definition 7, the latter is a collection of (precise) Markov semigroups $\mathbf{P}=(P_{t})_{t\geq 0}$, each associated with a Markov process $(X_{t})_{t\geq 0}$. In turn, every such Markov semigroup $\mathbf{P}$ encompasses a plausible invariant measure and a plausible transition probability. The scholar, then, can express their ambiguity around both, by letting the Markov processes $(X_{t})_{t\geq 0}$ associated with the elements $\mathbf{P}$ of $\mathbf{Q}$ have different invariant measures and transition probabilities.
As we shall see in Sections 3 and 4, taking an (imprecise Markov) semigroup route allows us to explicitly account for the geometry of the state space $E$ when we study the ergodic behavior of $\overline{\mathbf{P}}$ and $\underline{\mathbf{P}}$ (a perspective that is new to the literature), and to work in continuous time (whereas the majority of the results in the literature pertain the discrete time case). The price we pay for these gains is the need for a compact set of semigroups of linear operators. This is because, by Lemma 6, $\overline{\mathbf{P}},\underline{\mathbf{P}}\in\mathbf{Q}$, and so we can use techniques developed for (precise) Markov semigroups to study the ergodic behavior of $\overline{\mathbf{P}}$ and $\underline{\mathbf{P}}$, and in turn of all the elements of $\mathbf{Q}$. In the future, we plan to extend our approach involving the geometry of the state space $E$ to the frameworks introduced e.g. in Criens and Niemann (2024); Denk et al. (2018, 2020); Erreygers (2023); Feng and Zhao (2021); Kühn (2021); Nendel (2021), [^7] In those papers, the curvature of the state space $E$ is not considered by the authors when deriving their results. where nonlinear Markov semigroups (and in particular Markov semigroups associated with nonlinear Markov processes) are studied. This would allow us to retain the advantages of the semigroup approach, but without the restrictions coming from enforcing compactness of the imprecise Markov semigroup $\mathbf{Q}$, and linearity of its elements.
3. Ergodicity in the Riemannian Manifold Case
In this section, we begin our inspection on how the geometry of the state space $E$ informs us around the ergodicity of an imprecise Markov semigroup $\mathbf{Q}$. Our journey starts by letting $E=\mathbb{R}^{n}$, or $E=M$, a complete (with respect to the Riemannian distance) Riemannian manifold (Berger, 2003). [^8] For a primer on Riemannian geometry, we refer the reader to Bakry et al. (2014, Appendix C.3)
Loosely, a complete Riemannian manifold $M$ can be thought of as a “smooth” subset of $\mathbb{R}^{n}$, equipped with a complete metric that depends in some sense on the elements $x\in M$. A visual representation of a Riemannian manifold in $\mathbb{R}^{3}$ is given in Figure 3. In this context, the carré du champ and the iterated carré du champ operators associated with a Markov semigroup are easily computed (see B.1.1).
Theorem 8 (Limiting Behavior of $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$, $E$ Euclidean space or Riemannian manifold).
Let $E$ be $\mathbb{R}^{n}$ or a smooth complete connected manifold. Let $\mathcal{A}_{0}$ be the class of smooth ($\mathcal{C}^{\infty}$) compactly supported functionals $f$ on EE. [^9] Recall that a smooth function on a closed and bounded set (and hence, in this framework, compact by Heine-Borel) is itself bounded. Fix any $\tilde{f}\in B(E)$, and assume that we can define the partial order $\preceq_{\tilde{f}}^{\text{part}}$ or the total order $\preceq_{\tilde{f}}^{\text{tot}}$ that we introduced in Section 2. Let $\mathbf{Q}$ be an imprecise Markov semigroup. Denote by $\overline{\mathbf{P}}$ and $\underline{\mathbf{P}}$ the greatest and least elements of $\mathbf{Q}$, [^10] With respect to $\preceq^{\text{part}}_{\tilde{f}}$ or $\preceq^{\text{tot}}_{\tilde{f}}$. respectively, with invariant measures $\overline{\mu}$ and $\underline{\mu}$, and associated infinitesimal generators $\overline{L}$ and $\underline{L}$, respectively. Suppose that the following three conditions hold
  • (i)
    $\underline{L}$ and $\overline{L}$ are elliptic diffusion operators, symmetric with respect to $\underline{\mu}$ and $\overline{\mu}$, [^11] The symmetric assumption tell us that the law of the Markov processes associated with $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$ are “reversible in time”. respectively;
  • (ii)
    $\underline{L}$ and $\overline{L}$ satisfy the Bakry-Émery curvature conditions $CD(\rho_{\underline{L}},\infty)$ and $CD(\rho_{\overline{L}},\infty)$, respectively, for some $\rho_{\underline{L}},\rho_{\overline{L}}\in\mathbb{R}$;
  • (iii)
    The carré du champ algebra for both the carré du champ operators $\underline{\Gamma}$ and $\overline{\Gamma}$ of the Markov generators $\underline{L}$ and $\overline{L}$, respectively, is $\mathcal{A}_{0}$.
Then, we have that
$\forall f\in\mathbb{L}^{2}(\underline{\mu}),\quad\lim_{t\rightarrow\infty}\underline{P}_{t}f=\mathbb{E}_{\underline{\mu}}\left[f\right]=\int_{E}f\text{d}\underline{\mu},\quad\underline{\mu}\text{-a.e.}$
and
$\forall f\in\mathbb{L}^{2}(\overline{\mu}),\quad\lim_{t\rightarrow\infty}\overline{P}_{t}f=\mathbb{E}_{\overline{\mu}}\left[f\right]=\int_{E}f\text{d}\overline{\mu},\quad\overline{\mu}\text{-a.e.}$
in $\mathbb{L}^{2}(\underline{\mu})$ and in $\mathbb{L}^{2}(\overline{\mu})$, respectively. In addition, if the invariant measure for $\underline{\mathbf{P}}$ is equal to that of $\overline{\mathbf{P}}$, and if we denote it by $\mu$, then
$\forall f\in\mathbb{L}^{2}(\mu),\quad\lim_{t\rightarrow\infty}\overline{P}_{t}f=\lim_{t\rightarrow\infty}\underline{P}_{t}f=\mathbb{E}_{\mu}\left[f\right]=\int_{E}f\text{d}\mu,\quad\mu\text{-a.e.}$ (2)
in $\mathbb{L}^{2}(\mu)$.
If $E$ is a complete Riemannian manifold, the Bakry-Émery curvature condition $CD(\rho_{\underline{L}},\infty)$ in Theorem 8.(ii) is related to the notion of Ricci tensor at every point of $E$, and in turn to the Ricci curvature of the manifold $E$ (for more, see B.1.1). [^12] A similar argument holds for the infinitesimal generator $\overline{L}$ associated with Markov semigroup $\overline{\mathbf{P}}$. We see, then, how the geometry of the state space $E$ influences the ergodic behavior of an imprecise Markov semigroup.
Figure 3. In this figure, $M$ is a Riemannian manifold in $\mathbb{R}^{3}$, $x\in M$ is an element of the manifold, and $T_{x}M$ is the tangent space of $M$ at $x$ (the set of all tangent vectors to all smooth curves on $M$ passing through $x$). $\gamma(t)$ is a $\mathcal{C}^{1}$ curve passing through $x$, and $v\equiv v(x)=(v^{i}(x))_{i\in\{1,2,3\}}$ is a vector field.
Proof of Theorem 8.
We begin by pointing out that, by Lemma 6, $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$ are (precise) Markov semigroups, and so the infinitesimal generators $\underline{L}$ and $\overline{L}$ are well-defined.
We first focus on the greatest element $\overline{\mathbf{P}}$ of $\mathbf{Q}$. Recall that the variance $\text{Var}_{\overline{\mu}}(f)$ of function $f\in\mathbb{L}^{2}(\overline{\mu})$ is given by
$\text{Var}_{\overline{\mu}}(f)\coloneqq\int_{E}f^{2}\text{d}\overline{\mu}-\left(\int_{E}f\text{d}\overline{\mu}\right)^{2},$ (3)
and that the Dirichlet form $\overline{\mathcal{E}}(f)$ at $f\in\mathcal{A}_{0}$ of the carré du champ operator $\overline{\Gamma}$ of the Markov generator $\overline{L}$ is given by
$\overline{\mathcal{E}}(f)\coloneqq\int_{E}\overline{\Gamma}(f)\text{d}\overline{\mu}.$
Pick any $f\in\mathcal{A}_{0}$. Given our hypotheses, by Bakry et al. (2014, Theorem 3.2.3), we have that for every $t\geq 0$, the following gradient bound holds
$\overline{\Gamma}\left(\overline{P}_{t}f\right)\leq e^{-2t\rho_{\overline{L}}}\overline{P}_{t}\left(\overline{\Gamma}(f)\right).$ (4)
Then, by Bakry et al. (2014, Theorem 4.7.2), equation (4) implies that the following local Poincaré inequality holds for all $t\geq 0$,
$\overline{P}_{t}(f^{2})-\left(\overline{P}_{t}f\right)^{2}\leq\frac{1-e^{-2t\rho_{\overline{L}}}}{\rho_{\overline{L}}}\overline{P}_{t}\left(\overline{\Gamma}(f)\right).$ (5)
Following van Handel (2016, Problem 2.7), taking the limit as $t\rightarrow\infty$ on both sides of (5) yields the following Poincaré inequality
$\text{Var}_{\overline{\mu}}(f)\leq\frac{1}{\rho_{\overline{L}}}\overline{\mathcal{E}}(f).$ (6)
In addition, recall that for the Dirichlet form the Dirichlet form $\overline{\mathcal{E}}(f)$ we have that
$\overline{\mathcal{E}}(f)=-\int_{E}f\overline{L}f\text{d}\overline{\mu}.$ (7)
Combining (7) with (6), we have that if function $f$ is such that $\overline{L}f=0$, then $\overline{\mathcal{E}}(f)=0$. In turn, this implies that $\text{Var}_{\overline{\mu}}(f)=0$, and so the function $f$ must be constant $\overline{\mu}$-a.e. Notice now that, given that $\overline{L}$ is an elliptic operator, $\mathcal{A}_{0}$ is dense in $\mathcal{D}(\overline{L})$ in the topology induced by the infinitesimal generator domain norm $\|\cdot\|_{\mathcal{D}(\overline{L})}$. [^13] This norm is recalled in Definition 18.(DMT9) in B.2. Thanks to this, (3)-(7) hold for all $f\in\mathcal{D}(\overline{L})$ by an approximation argument. This implies that every $f\in\mathcal{D}(\overline{L})$ such that $\overline{L}f=0$ is constant, and so that $\overline{L}$ is ergodic according to Bakry et al. (2014, Definition 3.1.11). Then, since $\overline{\mu}$ is finite, by Bakry et al. (2014, Proposition 3.1.13) we have that
$\forall f\in\mathbb{L}^{2}(\overline{\mu}),\quad\lim_{t\rightarrow\infty}\overline{P}_{t}f=\int_{E}f\text{d}\overline{\mu},\quad\overline{\mu}\text{-a.e.}$
in $\mathbb{L}^{2}(\overline{\mu})$. This concludes the first part of the proof. The second part, that is, the one pertaining the least element $\underline{\mathbf{P}}$ of $\mathbf{Q}$, is analogous. The last statement of the theorem is immediate once we let $\overline{\mu}=\underline{\mu}=\mu$. ∎
Remark 1.
Let us add a technical remark to Theorem 8. If we do not know whether the (reversible invariant) measures $\overline{\mu}$ and $\underline{\mu}$ are finite, then in the statement of condition (ii) of Theorem 8 we must require that $\rho_{\underline{L}},\rho_{\overline{L}}>0$ (Bakry et al., 2014, Theorem 3.2.7), so that (4) holds.
Examples of semigroups for which the conditions (i)-(iii) of Theorem 8 are met are the Ornstein-Uhlenbeck and Laguerre semigroups (whose Markov generators satisfy the Bakry-Émery curvature condition with $\rho>0$; in particular, $\rho=1$ for the former, and $\rho=1/2$ for the latter), and the Brownian (Euclidean) semigroup (whose Markov generator satisfies the $CD(\rho,\infty)$ condition with $\rho=0$). In addition, an example of two Markov semigroups having the same invariant measure (as required for (2) to hold) are the Brownian motion with quadratic potential, and the Ornstein-Uhlenbeck semigroup with volatility parameter $\sigma=\sqrt{2}$. Indeed, for both, the invariant measure is a Gaussian with mean 0, and variance the inverse of the rate of mean reversion.
The following is an important corollary that subsumes the ergodic behavior of the elements of $\mathbf{Q}$ when applied to $\tilde{f}\in B(E)$.
Corollary 8.1 (Ergodicity of Imprecise Markov Semigroup $\mathbf{Q}$, $E$ Euclidean space or Riemannian manifold).
Retain the assumptions of Theorem 8, and let $\tilde{f}\in B(E)$ be the same functional that we fixed in Theorem 8. Then, for all $\mathbf{P}=(P_{t})_{t\geq 0}\in\mathbf{Q}$, we have that
$\lim_{t\rightarrow\infty}P_{t}\tilde{f}\geq\int_{E}\tilde{f}\text{d}\underline{\mu},\quad\underline{\mu}\text{-a.e.}$ (8)
in $\mathbb{L}^{2}(\underline{\mu})$, and
$\lim_{t\rightarrow\infty}P_{t}\tilde{f}\leq\int_{E}\tilde{f}\text{d}\overline{\mu},\quad\overline{\mu}\text{-a.e.}$ (9)
in $\mathbb{L}^{2}(\overline{\mu})$. In addition, if the invariant measure for $\underline{\mathbf{P}}$ is equal to that of $\overline{\mathbf{P}}$, and if we denote it by $\mu$, then
$\lim_{t\rightarrow\infty}P_{t}\tilde{f}=\int_{E}\tilde{f}\text{d}{\mu},\quad{\mu}\text{-a.e.}$ (10)
in $\mathbb{L}^{2}(\mu)$.
Proof.
[…]
Let us add some interesting comments on Corollary 8.1. In (10), we see how, if the initial distribution – i.e. the distribution of $X_{0}$ – is known and equal to $\mu$, then the ambiguity that the agent faces around the correct transition probability of the process from $X_{0}$ to $X_{t}$, vanishes as $t\rightarrow\infty$. We recover the initial (reversible invariant) distribution $\mu$, according to which we compute the expectations. As the classical ergodic adage goes, “in the limit, time average ($\lim_{t\rightarrow\infty}\underline{P}_{t}\tilde{f}=\lim_{t\rightarrow\infty}\overline{P}_{t}\tilde{f}$) equals space average ($\int_{E}\tilde{f}\text{d}\mu$)”. We say that an imprecise Markov semigroup $\mathbf{Q}$ is $\tilde{f}-$lower ergodic if it satisfies (8), $\tilde{f}-$upper ergodic if it satisfies (9), and $\tilde{f}-$ergodic if it satisfies (10).
As we pointed out in Section 2, there are two main differences with respect to the existing literatures on imprecise Markov processes and ergodicity for imprecise probabilities. We are the first to explicitly exploit the geometry of the state space $E$ via the Bakry-Émery curvature condition to derive our results. In addition, in Theorem 8 and Corollary 8.1 we represent the scholar’s ambiguity about the invariant measure and the transition probability via an imprecise Markov semigroup $\mathbf{Q}$. We briefly mention that our approach differs slightly from the imprecise ergodic literature (see e.g. Caprio and Gong (2023); Caprio and Mukherjee (2023); Cerreia-Vioglio et al. (2015)), where ambiguity around distribution $\mu$ is typically represented via a Choquet approach involving a superadditive Choquet capacity (Choquet, 1954).
We conclude this section with a corollary that further informs us on the geometrical properties of $E$ in the ergodic regime of $\mathbf{Q}$, when $E$ is a complete Riemannian manifold. Let us denote by $E^{n}_{K}$ the complete $n$-dimensional simply connected space of constant sectional curvature $K\in\mathbb{R}$, [^14] The Ricci curvature of $E^{n}_{K}$ is constant and equal to (n1)K(n-1)K. and by $B(p,r)$ the ball of radius $r$ at point $p$, defined with respect to the Riemannian distance function.
Corollary 8.2 (Bishop–Gromov Inequality).
Let $E\subseteq\mathbb{R}^{n}$ be a complete Riemannian manifold. Retain the first part of Theorem 8, and assume that condition (ii) holds. Then,
$\text{Vol}\left[B(p,r)\right]\leq\text{Vol}\left[B(p_{K},r)\right],$
for all $p\in E$, all $p_{K}\in E^{n}_{K}$, all $r\in(0,\infty)$, and some $K$ that depends on $\rho_{\underline{L}}$ and $\rho_{\overline{L}}$. Here, $\text{Vol}[\cdot]$ denotes the volume operator.
Proof.
Given our assumptions, we know that $CD(\rho_{\underline{L}},\infty)$ and $CD(\rho_{\overline{L}},\infty)$ both hold. Let us define $\rho_{\star}\coloneqq\min\{\rho_{\underline{L}},\rho_{\overline{L}}\}$, and put $\rho_{\star}=(n-1)K$, so that $K=\rho_{\star}/(n-1)$. The Bakry-Émery curvature assumptions imply that the Ricci tensor at every point of $E$ is bounded from below by $\rho_{\star}$. [^15] This is discussed in B.1.1. In turn, we have that the Ricci curvature Ric of $E$ is lower bounded by (n1)K(n-1)K, that is, $\text{Ric}\geq(n-1)K$. The result then follows from Bishop (1963). ∎
Corollary 8.2 tells us that – under the Bakry-Émery curvature condition (ii) of Theorem 8 – the volume of a ball around any element of the manifold $E$ is upper bounded by the volume of a ball around any element of $E^{n}_{K}$ having the same radius. Further comparison results like Corollary 8.2 that hold in the ergodic regime of $\mathbf{Q}$ will be the subject of future study.
4. Ergodicity in the General Case
In this section, our goal is to extend Theorem 8 and Corollary 8.1 to the case where $(E,\mathcal{F})$ is an arbitrary good measurable space. To do so, we need to introduce the concept of an Imprecise Diffusion Markov Tuple, the imprecise version of a diffusion Markov triple. The latter is recalled in B.2.
Definition 9 (Imprecise Diffusion Markov (5-)Tuple).
Fix any $\tilde{f}\in B(E)$, and assume that we can define the partial order $\preceq_{\tilde{f}}^{\text{part}}$ or the total order $\preceq_{\tilde{f}}^{\text{tot}}$ that we introduced in Section 2. An imprecise diffusion Markov tuple $(E,\overline{\mu},\underline{\mu},\overline{\Gamma},\underline{\Gamma})$ is a 5-tuple such that the following holds.
  • $(E,\mathcal{F})$ is a good measurable space.
  • $\overline{\mu}$ and $\underline{\mu}$ are measures on $\mathcal{F}$.
  • $\mathcal{A}^{\overline{\mu}}_{0}$ and $\mathcal{A}^{\underline{\mu}}_{0}$ are vector spaces of bounded measurable functionals dense in all $\mathbb{L}^{p}(\overline{\mu})$ and all $\mathbb{L}^{p}(\underline{\mu})$ spaces, $1\leq p<\infty$, respectively, and stable under products.
  • $(E,\overline{\mu},\overline{\Gamma})$ and $(E,\underline{\mu},\underline{\Gamma})$ are diffusion Markov triples having associated algebras $\mathcal{A}^{\overline{\mu}}_{0}$ and $\mathcal{A}^{\underline{\mu}}_{0}$, respectively, and carré du champs operators $\overline{\Gamma}$ and $\underline{\Gamma}$, respectively.
  • The Markov semigroups $\overline{\mathbf{P}}=(\overline{P}_{t})_{t\geq 0}$ and $\underline{\mathbf{P}}=(\underline{P}_{t})_{t\geq 0}$ associated with $\overline{\Gamma}$ and $\underline{\Gamma}$ are the greatest and least elements of an imprecise Markov semigroup $\mathbf{Q}$, [^16] With respect to $\preceq^{\text{part}}_{\tilde{f}}$ or $\preceq^{\text{tot}}_{\tilde{f}}$. respectively.
We are now ready for the main result of this section. It hinges on the concepts of adjoint operator, self-adjointness, essential self-adjointness (ESA), extended algebra $\mathcal{A}$ of the algebra $\mathcal{A}_{0}$ associated with a diffusion Markov triple, connexity, and weak hypo-ellipticity. They are brushed off in B.2. As its proof technique is similar to that of Theorem 8 (despite exploiting different, more general results), we defer it to A.
Theorem 10 (Limiting Behavior of $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$, General).
Fix any $\tilde{f}\in B(E)$, and assume that we can define the partial order $\preceq_{\tilde{f}}^{\text{part}}$ or the total order $\preceq_{\tilde{f}}^{\text{tot}}$ that we introduced in Section 2. Let $(E,\overline{\mu},\underline{\mu},\overline{\Gamma},\underline{\Gamma})$ be an imprecise diffusion Markov tuple, and denote by $\mathcal{A}^{\overline{\mu}}$ and $\mathcal{A}^{\underline{\mu}}$ the extended algebras of $\mathcal{A}_{0}^{\overline{\mu}}$ and $\mathcal{A}_{0}^{\underline{\mu}}$ associated with $(E,\overline{\mu},\overline{\Gamma})$ and $(E,\underline{\mu},\underline{\Gamma})$, respectively. If the connexity and weak hypo-ellipticity assumptions hold for both $(E,\overline{\mu},\overline{\Gamma})$ and $(E,\underline{\mu},\underline{\Gamma})$, and if the Bakry-Émery curvature conditions $CD(\rho_{\overline{L}},\infty)$ and $CD(\rho_{\underline{L}},\infty)$ are met by both the infinitesimal generators $\overline{L}$ and $\underline{L}$ associated with $\overline{\Gamma}$ and $\underline{\Gamma}$, respectively, for some $\rho_{\overline{L}},\rho_{\underline{L}}\in\mathbb{R}$, then the following hold
$\forall f\in\mathbb{L}^{2}(\underline{\mu}),\quad\lim_{t\rightarrow\infty}\underline{P}_{t}f=\mathbb{E}_{\underline{\mu}}[f]=\int_{E}f\text{d}\underline{\mu}\quad\text{in }\mathbb{L}^{2}(\underline{\mu}),\quad\underline{\mu}\text{-a.e.}$
and
$\forall f\in\mathbb{L}^{2}(\overline{\mu}),\quad\lim_{t\rightarrow\infty}\overline{P}_{t}f=\mathbb{E}_{\overline{\mu}}[f]=\int_{E}f\text{d}\overline{\mu}\quad\text{in }\mathbb{L}^{2}(\overline{\mu}),\quad\overline{\mu}\text{-a.e.}$
In addition, if $\overline{\mu}=\underline{\mu}=\mu$, then
$\forall f\in\mathbb{L}^{2}(\mu),\quad\lim_{t\rightarrow\infty}\overline{P}_{t}f=\lim_{t\rightarrow\infty}\underline{P}_{t}f=\mathbb{E}_{\mu}[f]=\int_{E}f\text{d}\mu\quad\text{in }\mathbb{L}^{2}(\mu),\quad\mu\text{-a.e.}$
The connexity and weak hypo-ellipticity assumptions allow to “import” some of the geometric structure of the Riemannian manifolds considered in Section 3 to the more general diffusion Markov triple setting, and to relax the requirement in Theorem 8 that $\underline{L}$ and $\overline{L}$ are elliptic operators. This is expanded upon in B.2.
Similarly to what we pointed out in Remark 1, if we do not know whether the (reversible invariant) measures $\overline{\mu}$ and $\underline{\mu}$ are finite, then in the statement of Theorem 10 we must require that $\rho_{\overline{L}},\rho_{\underline{L}}>0$ (Bakry et al., 2014, Theorem 3.3.23). [^17] So that (11) holds.
Similarly to Corollary 8.1, the following is an important result that subsumes the ergodic behavior of the elements of $\mathbf{Q}$ when applied to $\tilde{f}\in B(E)$.
Corollary 10.1 (Ergodicity of Imprecise Markov Semigroup $\mathbf{Q}$, General).
Retain the assumptions of Theorem 10, and let $\tilde{f}\in B(E)$ be the same functional that we fixed in Theorem 10. Then, for all $\mathbf{P}=(P_{t})_{t\geq 0}\in\mathbf{Q}$, we have that
$\lim_{t\rightarrow\infty}P_{t}\tilde{f}\geq\int_{E}\tilde{f}\text{d}\underline{\mu},\quad\underline{\mu}\text{-a.e.}$
in $\mathbb{L}^{2}(\underline{\mu})$, and
$\lim_{t\rightarrow\infty}P_{t}\tilde{f}\leq\int_{E}\tilde{f}\text{d}\overline{\mu},\quad\overline{\mu}\text{-a.e.}$
in $\mathbb{L}^{2}(\overline{\mu})$. In addition, if the invariant measure for $\underline{\mathbf{P}}$ is equal to that of $\overline{\mathbf{P}}$, and if we denote it by $\mu$, then
$\lim_{t\rightarrow\infty}P_{t}\tilde{f}=\int_{E}\tilde{f}\text{d}{\mu},\quad{\mu}\text{-a.e.}$
in $\mathbb{L}^{2}(\mu)$.
Proof.
[…]
5. Conclusion
In this work, we introduced the concept of an imprecise Markov semigroup, and studied its $\tilde{f}-$ergodic behavior, for some $\tilde{f}\in B(E)$, both when the state space $E$ is a Euclidean space or a Riemannian manifold, and when $E$ is allowed to be an arbitrary good measurable space.
In the future, we plan to forego the assumption that $\mathbf{Q}$ is compact, and work directly with nonlinear Markov semigroups (Kühn, 2021; Feng and Zhao, 2021). This will require a considerable effort, since a new theory involving the Bakry-Émery curvature condition associated with nonlinear Markov semigroups will need to be developed.
Before this (ambitious) goal, we plan to apply our findings to machine learning and computer vision, as we discussed in the Introduction, and to achieve robustness in reinforcement learning, where Markov processes are used for planning in autonomous agents. We also aim to extend to the imprecise case the types of inequalities put forth in Bakry et al. (2014, Chapters 6-9). In particular, we are interested in deriving imprecise optimal transport inequalities, similarly to (Caprio, 2024; Lorenzini et al., 2024), with the intent of applying them to distribution shift problems in machine learning and computer vision (Lin et al., 2023).
Acknowledgements
We would like to acknowledge partial funding by the Army Research Office (ARO MURI W911NF2010080). We wish to express our gratitude to Mira Gordin for insightful discussions, particularly on how the gradient bound for a diffusion operator relates to Poincaré inequalities, and to Simone Cerreia-Vioglio and Paolo Perrone for their help in finding the suitable topology compatible with the partial order $\preceq^{\text{part}}_{\tilde{f}}$, and in shaping up the discussion after Lemma 6. We are also extremely grateful to Enrique Miranda for his suggestion to look into Krak et al. (2017, Example 6.2), to Andrea Aveni for his comments on an initial version of this work, and to Malena Español for her invitation to the Institute for Advanced Study, where the idea of the present paper germinated.
Appendix A Proof of Theorem 10
We begin by pointing out that, by Lemma 6 and Definition 9, $\underline{\mathbf{P}}$ and $\overline{\mathbf{P}}$ are (precise) Markov semigroups, and so the operators $\underline{\Gamma}$, $\overline{\Gamma}$, $\underline{L}$ and $\overline{L}$ are well-defined.
We first focus on the diffusion Markov triple $(E,\overline{\mu},\overline{\Gamma})$. By Bakry et al. (2014, Proposition 3.3.11) and the fact that $\overline{\mathbf{P}}=(\overline{P}_{t})_{t\geq 0}$ is a (precise) Markov semigroup, we have that $\overline{L}$ is essentially self-adjoint. Pick any $f\in\mathcal{A}_{0}^{\overline{\mu}}$. The fact that $\overline{L}$ satisfies ESA, together with our assumption that it satisfies the curvature condition $CD(\rho_{\overline{L}},\infty)$, for some $\rho_{\overline{L}}\in\mathbb{R}$, implies – by Bakry et al. (2014, Corollary 3.3.19) – that for all $t\geq 0$, the following gradient bound holds
$\overline{\Gamma}\left(\overline{P}_{t}f\right)\leq e^{-2t\rho_{\overline{L}}}\overline{P}_{t}\left(\overline{\Gamma}(f)\right).$ (11)
Then, by Bakry et al. (2014, Theorem 4.7.2), equation (11) implies that the following local Poincaré inequality holds for all $t\geq 0$,
$\overline{P}_{t}(f^{2})-\left(\overline{P}_{t}f\right)^{2}\leq\frac{1-e^{-2t\rho_{\overline{L}}}}{\rho_{\overline{L}}}\overline{P}_{t}\left(\overline{\Gamma}(f)\right).$ (12)
Following van Handel (2016, Problem 2.7), taking the limit as $t\rightarrow\infty$ on both sides of (12) yields the following Poincaré inequality
$\text{Var}_{\overline{\mu}}(f)\leq\frac{1}{\rho_{\overline{L}}}\overline{\mathcal{E}}(f).$ (13)
In addition, we know that
$\overline{\mathcal{E}}(f)=-\int_{E}f\overline{L}f\text{d}\overline{\mu}.$ (14)
Combining (14) with (13), we have that if function $f$ is such that $\overline{L}f=0$, then $\overline{\mathcal{E}}(f)=0$. In turn, this implies that $\text{Var}_{\overline{\mu}}(f)=0$, and so the function $f$ must be constant $\overline{\mu}$-a.e. Notice now that, since $\overline{L}$ satisfies the ESA property, $\mathcal{A}_{0}^{\overline{\mu}}$ is dense in $\mathcal{D}(\overline{\mathcal{E}})$ in the topology induced by the Dirichlet norm $\|\cdot\|_{\overline{\mathcal{E}}}$. [^18] This norm is recalled in Definition 18.(DMT4) in B.2. Thanks to this, (11)-(14) hold for all $f\in\mathcal{D}(\overline{\mathcal{E}})$ by an approximation argument. This implies that every $f\in\mathcal{D}(\overline{\mathcal{E}})$ such that $\overline{\Gamma}(f)=0$ is constant, and so that $\overline{L}$ is ergodic according to Bakry et al. (2014, Definition 3.1.11). Then, since $\overline{\mu}$ is finite, by Bakry et al. (2014, Proposition 3.1.13) we have that
$\forall f\in\mathbb{L}^{2}(\overline{\mu}),\quad\lim_{t\rightarrow\infty}\overline{P}_{t}f=\int_{E}f\text{d}\overline{\mu},\quad\overline{\mu}\text{-a.e.}$
in $\mathbb{L}^{2}(\overline{\mu})$. This concludes the first part of the proof. The second part, that is, the one pertaining the least element $\underline{\mathbf{P}}$ of $\mathbf{Q}$, is analogous. The last statement of the theorem is immediate once we let $\overline{\mu}=\underline{\mu}=\mu$. ∎
Appendix B Background Notions
In an effort to keep the paper self-contained, we present in this section the preliminary concepts needed to appreciate the main results in our work.
B.1. Background on Markov Semigroups
Let us briefly recall what a (precise) Markov process is; to do so, we borrow the notation and terminology from Bakry et al. (2014, Chapter 1.1). Let $(X_{t})_{t\geq 0}$ be a measurable process on a probability space $(\Omega,\Sigma,\mathbb{P})$, such that $X_{t}\equiv X_{t}(\omega)$ is an element of the good measurable state space $E$ from Definition 1, for all $t\geq 0$ and all $\omega\in\Omega$. Denote by $\mathcal{F}_{t}\coloneqq\sigma(X_{u}:u\leq t)$, $t\geq 0$, the natural filtration of $(X_{t})_{t\geq 0}$.
The Markov property indicates that for $t>s$, the law of $X_{t}$ given $\mathcal{F}_{s}$ is the law of $X_{t}$ given $X_{s}$, as well as the law of $X_{t-s}$ given $X_{0}$, the latter property reflecting the fact that the Markov process is time homogeneous, which is the only case that we consider in the present work. As a consequence, throughout the paper we focus our attention on the law of $X_{t}$ given $X_{0}$, for a generic $t\geq 0$.
For any $t\geq 0$, we can describe the distribution of $X_{t}$ starting from $X_{0}=x$ via a probability kernel $p_{t}(x,A)$, that is, a function $p_{t}:E\times\mathcal{F}\rightarrow[0,1]$ such that (i) $p_{t}(x,\cdot)$ is a probability measure, for all $x\in E$, and (ii) $x\mapsto p_{t}(x,A)$ is measurable, for every measurable set $A\in\mathcal{F}$.
We now introduce one of the main concepts that we use in the present paper, namely a Markov semigroup. We follow the presentation in (Bakry et al., 2014, Chapter 1.2). Let $B(E)$ denote the set of bounded measurable functionals on $(E,\mathcal{F})$. A Markov semigroup $\mathbf{P}=(P_{t})_{t\geq 0}$ is a family of operators that satisfies the following properties.
  • (i)
    For every $t\geq 0$, $P_{t}:B(E)\rightarrow B(E)$ is a linear operator (linearity).
  • (ii)
    $P_{0}=\text{Id}$, the identity operator (initial condition).
  • (iii)
    $P_{t}\mathbbm{1}=\mathbbm{1}$, where $\mathbbm{1}$ is the constant function equal to $1$ (mass conservation). [^19] In general, we write $P_{t}(f)\equiv P_{t}f$, for all $f\in B(E)$, for notational convenience.
  • (iv)
    For every $t\geq 0$, $f\geq 0\implies P_{t}f\geq 0$ (positivity preserving).
  • (v)
    For every $t,s\geq 0$, $P_{t+s}=P_{t}\circ P_{s}$ (semigroup property).
One extra condition – continuity at $t=0$ – is needed to fully define a Markov semigroup $\mathbf{P}=(P_{t})_{t\geq 0}$. To present it though, we first need to introduce the concept of invariant (or stationary) measure.
Let $\mathbf{P}=(P_{t})_{t\geq 0}$ be a family of operators satisfying (i)-(v). Then, a (positive) $\sigma$-finite measure $\mu$ on $(E,\mathcal{F})$ is called invariant for $\mathbf{P}$ if for every bounded positive measurable functional $f$ on $E$, and every $t\geq 0$,
$\int_{E}P_{t}f\text{d}\mu=\int_{E}f\text{d}\mu.$ (15)
Notice that, as a consequence of (iii), (iv), and Jensen’s inequality, we have that if a measure $\mu$ is invariant for $\mathbf{P}$, then, for every $t\geq 0$, $P_{t}$ is a contraction on the bounded functions in $\mathbb{L}^{p}(\mu)$, for any $1\leq p\leq\infty$.
We are now ready for the complete definition.
Definition 11 (Markov Semigroup).
Let $(E,\mathcal{F})$ be a good measurable space. A family $\mathbf{P}=(P_{t})_{t\geq 0}$ of operators $P_{t}$ is called a Markov semigroup if it satisfies (i)-(v), there exists an invariant measure $\mu$ for $\mathbf{P}$, and the following extra condition is satisfied
  • (vi)
    For every $f\in\mathbb{L}^{2}(\mu)$, $\lim_{t\rightarrow 0}P_{t}f=f$ in $\mathbb{L}^{2}(\mu)$ (continuity at $t=0$).
Notice that, since $(P_{t})_{t\geq 0}$ is a contraction (as pointed out after equation (15)) and a semigroup (by condition (v)), property (vi) expresses that $t\mapsto P_{t}f$ is continuous in $\mathbb{L}^{2}(\mu)$ on $\mathbb{R}_{+}$. In addition, we could have required convergence in $\mathbb{L}^{p}(\mu)$ for any $p\in[1,\infty)$, but we stick to $p=2$ for simplicity and to align with Bakry et al. (2014). Before moving on, let us add a remark on invariant measures.
Remark 2 (Spotlight on Invariant Measures).
In general, an invariant measure $\mu$ for a Markov semigroup $\mathbf{P}$ need not be a probability measure. [^20] If it is a finite measure, it is customary to normalize it to a probability measure. It is then usually unique (Bakry et al., 2014, Page 11). In addition, it need not always exist; necessary and sufficient conditions for its existence may be found in (Ito, 1964; Horowitz, 1974). That being said, most semigroups of interest do have an invariant measure (Bakry et al., 2014, Section 1.2.1). A way to derive $\mu$ as a weak limit of a “reasonable” initial probability measure $\mu_{0}$ is inspected in Bakry et al. (2014, Page 10). A typical example of a Markov semigroup having an invariant measure is the heat or Brownian semigroup on $\mathbb{R}^{n}$, whose invariant measure is (up to multiplication by a constant) the Lebesgue measure. [^21] The invariant measure is in general only defined up to a multiplicative constant.
As pointed out in Bakry et al. (2014, Section 1.2.1), when the invariant measure $\mu$ is a probability measure, it has an immediate interpretation. If the Markov process $(X_{t})_{t\geq 0}$ starts at time $t=0$ with initial distribution $\mu$, i.e. if $X_{0}\sim\mu$, then it keeps the same distribution at each time $t\geq 0$; indeed,
$\mathbb{E}[f(X_{t})]=\mathbb{E}[\mathbb{E}(f(X_{t})\mid X_{0})]=\mathbb{E}[P_{t}f(X_{0})]$ $=\int_{E}P_{t}f\text{d}\mu$
$=\int_{E}f\text{d}\mu=\mathbb{E}[f(X_{0})],$
for all $f\in B(E)$ and all $t\geq 0$. Functionals $f$, then, have to be understood as classes of functions for the $\mu$-a.e. equality, and equalities and inequalities such as $f\leq g$ are always understood to hold $\mu$-a.e.
We now continue our presentation of the concepts related to Markov semigroups that are crucial to derive our main results. An important property of a Markov semigroup is its symmetry (Bakry et al., 2014, Definition 1.6.1).
Definition 12 (Symmetric Markov Semigroup).
Let $\mathbf{P}=(P_{t})_{t\geq 0}$ be a Markov semigroup with (good) state space $(E,\mathcal{F})$ and invariant measure $\mu$. We say $\mathbf{P}$ is symmetric with respect to $\mu$ – or equivalently, that $\mu$ is reversible for $\mathbf{P}$ – if, for all functions $f,g\in\mathbb{L}^{2}(\mu)$ and all $t\geq 0$,
$\int_{E}fP_{t}g\text{d}\mu=\int_{E}gP_{t}f\text{d}\mu.$
The probabilistic interpretation of Definition 12 is straightforward. As pointed out in Bakry et al. (2014, Section 1.6.1), the name “reversible” refers to reversibility in time of the Markov process $(X_{t})_{t\geq 0}$ associated with the Markov semigroup $\mathbf{P}$ whenever the initial law $\mu$ is the invariant measure. Indeed, from (15), we know that if the process starts from the invariant distribution $\mu$, then it keeps the same distribution at each time $t$. Moreover, if the measure is reversible, and if the initial distribution of $X_{0}$ is $\mu$, then for any $t>0$ and any partition $0\leq t_{1}\leq\cdots\leq t_{k}\leq t$ of the time interval $[0,t]$, the law of $(X_{0},X_{t_{1}},\ldots,X_{t_{k}},X_{t})$ is the same as the law of $(X_{t},X_{t-t_{1}},\ldots,X_{t-t_{k}},X_{0})$. Hence, we can say that the law of the Markov process is “reversible in time”. [^22] For examples of symmetric Markov semigroups, we refer the reader to Bakry et al. (2014, Section 1.6.1, Appendices A.2, A.3).
We now need to introduce the concept of infinitesimal generator of a Markov semigroup $\mathbf{P}=(P_{t})_{t\geq 0}$ (Bakry et al., 2014, Definition 1.4.1). By Hille-Yosida theory (Hille and Phillips, 1996), [^23] A more modern reference for this result is (Bakry et al., 2014, Appendix A.1). there exists a dense linear subspace of $\mathbb{L}^{2}(\mu)$, called the domain $\mathcal{D}$ of the semigroup $\mathbf{P}$, on which the derivative at $t=0$ of $P_{t}$ exists in $\mathbb{L}^{2}(\mu)$.
Definition 13 (Infinitesimal Generator).
Let $\mathbf{P}=(P_{t})_{t\geq 0}$ be a Markov semigroup with (good) state space $(E,\mathcal{F})$ and invariant measure $\mu$. The infinitesimal generator $L$ of $\mathbf{P}$ in $\mathbb{L}^{2}(\mu)$ is a linear operator $L:\mathcal{D}\rightarrow\mathbb{L}^{2}(\mu)$,
$f\mapsto Lf\coloneqq\lim_{t\rightarrow 0}\frac{P_{t}f-f}{t}.$
The domain $\mathcal{D}$ is also called the domain of $L$, denoted by $\mathcal{D}(L)$, and depends on the underlying space $\mathbb{L}^{2}(\mu)$. Two comments are in place. First, as indicated by Bakry et al. (2014, Chapter 1.4), the couple $(L,\mathcal{D}(L))$ completely characterizes the Markov semigroup $\mathbf{P}$ acting on $\mathbb{L}^{2}(\mu)$. Second, the infinitesimal generator $L$ has a natural interpretation. As pointed out by Dan MacKinlay, $L$ is a kind of linearization of the local Markov transition kernel for the Markov process $(X_{t})_{t\geq 0}$ associated with the Markov semigroup $\mathbf{P}$; this is because $L$ is defined as a derivative.
The infinitesimal generator $L$ of a Markov semigroup $\mathbf{P}$ is sometimes referred to as the Markov generator (of $\mathbf{P}$). We can express the reversibility property of Definition 12 in terms of $L$ as
$\int_{E}fLg\text{d}\mu=\int_{E}gLf\text{d}\mu,\quad\forall f,g\in\mathcal{D}(L).$ (16)
The Markov generator is thus symmetric, however it is unbounded in $\mathbb{L}^{2}(\mu)$.
Next, we present the carré du champ operator, that measures how far an infinitesimal generator $L$ is from being a derivation (Ledoux, 2000). [^24] Here, “derivation” refers to the differential algebra meaning of the term (Lang, 2002).
Definition 14 (Carré du Champ).
Let $\mathbf{P}=(P_{t})_{t\geq 0}$ be a Markov semigroup with (good) state space $(E,\mathcal{F})$ and invariant measure $\mu$. Let $L$ be its infinitesimal generator with $\mathbb{L}^{2}(\mu)$-domain $\mathcal{D}(L)$. Let $\mathcal{A}\subseteq\mathcal{D}(L)$ be an algebra. Then, the bilinear map
$\Gamma:\mathcal{A}\times\mathcal{A}\rightarrow\mathcal{A},\quad(f,g)\mapsto\Gamma(f,g)\coloneqq\frac{1}{2}\left[L(fg)-fLg-gLf\right]$
is called the carré du champ operator of the Markov generator $L$.
The carré du champ operator is symmetric and is positive on $\mathcal{A}$, that is, $\Gamma(f,f)\geq 0$, for all $f\in\mathcal{A}$. [^25] This is a fundamental property of Markov semigroups (Bakry et al., 2014, Section 1.4.2). In this work, we write $\Gamma(f,f)\equiv\Gamma(f)$, $f\in\mathcal{A}$, to lighten the notation. In addition, notice that the definition of the carré du champ operator is subordinate to the algebra $\mathcal{A}$. For this reason, we call the latter the carré du champ algebra.
We now ask ourselves whether it is enough to know the Markov generator $L$ on smooth functionals on $E$ in order to describe the associated Markov semigroup $\mathbf{P}=(P_{t})_{t\geq 0}$. The answer is positive, provided $L$ is a diffusion operator (Bakry et al., 2014, Definition 1.11.1).
Definition 15 (Diffusion Operator).
An operator $L$, with carré du champ operator $\Gamma$, is said to be a diffusion operator if
$L\psi(f)=\psi^{\prime}(f)Lf+\psi^{\prime\prime}(f)\Gamma(f),$ (17)
for all $\psi:\mathbb{R}\rightarrow\mathbb{R}$ of class at least $\mathcal{C}^{2}$, and every suitably smooth functional $f$ on EE. [^26] Typically of class at least $\mathcal{C}^{2}$.
Together with the following (Bakry et al., 2014, Section 1.16.1), the notion of diffusion operator will allow us to link the probabilistic properties of a Markov semigroup to the geometry of the state space $E$.
Definition 16 (Iterated Carré du Champ Operator).
Let $\mathbf{P}=(P_{t})_{t\geq 0}$ be a symmetric Markov semigroup with (good) state space $(E,\mathcal{F})$ and invariant reversible measure $\mu$. Let $L$ be its infinitesimal generator with $\mathbb{L}^{2}(\mu)$-domain $\mathcal{D}(L)$, and denote by $\mathcal{A}$ the algebra on which the carré du champ operator $\Gamma$ of $L$ is well defined. The iterated carré du champ operator of the Markov generator $L$ is defined as
$\mathcal{A}\ni\Gamma_{2}(f,g)\coloneqq\frac{1}{2}\left[L\Gamma(fg)-\Gamma(f,Lg)-\Gamma(Lf,g)\right],$ (18)
for every pair $(f,g)\in\mathcal{A}\times\mathcal{A}$ such that the terms on the right hand side of (18) are well defined.
As for the carré du champ operator $\Gamma$, we write $\Gamma_{2}(f,f)\equiv\Gamma_{2}(f)$ to lighten the notation. Unlike $\Gamma$, $\Gamma_{2}$ is not always positive. As we will see more formally in B.1.1, $\Gamma_{2}$ is positive if and only if the state space $E$ has “positive curvature”. This high-level intuition, together with the fact that many differential inequalities can be expressed via $\Gamma_{2}$ and $\Gamma$, justifies the following definition (Bakry et al., 2014, Definition 1.16.1).
Definition 17 (Bakry-Émery Curvature).
Let $\mathbf{P}=(P_{t})_{t\geq 0}$ be a symmetric Markov semigroup with (good) state space $(E,\mathcal{F})$ and invariant reversible measure $\mu$. Let $L$ be its infinitesimal generator with $\mathbb{L}^{2}(\mu)$-domain $\mathcal{D}(L)$, and assume it is a diffusion operator. Denote by $\mathcal{A}$ the algebra on which the carré du champ operator $\Gamma$ of $L$ is well defined. [^27] E.g. the algebra of functionals of class at least $\mathcal{C}^{2}$ on $E$. Operator $L$ is said to satisfy the Bakry-Émery curvature condition $CD(\rho,n)$, for $\rho\in\mathbb{R}$ and $n\in[1,\infty]$, if for every functional $f$ in $\mathcal{A}$, the following holds $\mu$-a.e.
$\Gamma_{2}(f)\geq\rho\Gamma(f)+\frac{1}{n}(Lf)^{2}.$ (19)
Of course, equation (19) becomes $\Gamma_{2}(f)\geq\rho\Gamma(f)$ for $n=\infty$.
B.1.1. Riemannian Manifold
When $E=\mathbb{R}^{n}$, or $E=M$, a complete (with respect to the Riemannian distance) Riemannian manifold (see Figure 3), the carré du champ and the iterated carré du champ operators are easily computed.
[…]
$\Gamma_{2}(f)\geq\rho\Gamma(f)=\rho|\nabla f|^{2}\text{, }\forall f\in\mathcal{A}_{0}\iff\text{Ric}(\nabla f,\nabla f)\geq\rho|\nabla f|^{2},$
that is, the Bakry-Émery curvature condition $CD(\rho,\infty)$ holds if and only if the Ricci tensor at every point is bounded from below by $\rho$. In turn, $\Gamma_{2}$ is positive if and only if the Ricci curvature Ric of the manifold is positive. This makes more precise the idea we expressed in the previous section that $\Gamma_{2}$ is positive if and only if the state space $E$ has “positive curvature”.
We now introduce the concept of elliptic diffusion operator. First, notice that when $E$ is $\mathbb{R}^{n}$ or a Riemannian manifold $M$, we can rewrite condition (17) for $L$ to be a diffusion operator as
$Lf=\sum_{i,j\in\{1,\ldots,n\}}g^{ij}\partial^{2}_{ij}f+\sum_{i=1}^{n}b^{i}\partial_{i}f,\quad f\in\mathcal{A}_{0}.$ (20)
Here, $\mathfrak{g}\equiv\mathfrak{g}(x)=(g^{ij}(x))_{i,j\in\{1,\ldots,n\}}$ is a smooth ($\mathcal{C}^{\infty}$) $n\times n$ symmetric matrix-valued function of $x\in E$ (e.g. the co-metric of $M$ at $x$), and $b\equiv b(x)=(b^{i}(x))_{i=1}^{n}$ is a smooth ($\mathcal{C}^{\infty}$) $\mathbb{R}^{n}$-valued function of $x\in E$ (e.g. a vector field). In turn, this implies that we can write the carré du champ operator as $\Gamma(f,g)=\sum_{i,j\in\{1,\ldots,n\}}g^{ij}\partial_{i}f\partial_{j}g$, $f,g\in\mathcal{A}_{0}$. A diffusion operator as in (20) is said to be elliptic if
$\mathfrak{g}(V,V)=\sum_{i,j\in\{1,\ldots,n\}}g^{ij}(x)V_{i}V_{j}\geq 0\text{, }\forall V=(V_{i})_{i=1}^{n}\in\mathbb{R}^{n}$
and
$\mathfrak{g}(V,V)=0\implies V=0.$
Loosely, the fact that a diffusion operator $L$ is elliptic is important because it ensures that any solution on $0,s]\times\mathcal{O}$ of the heat equation $\partial_{t}u=Lu$, $u(0,x)=u_{0}$, where $s\in\mathbb{R}_{>0}$ and $\mathcal{O}\subseteq\mathbb{R}^{n}$ is open in the Euclidean topology, is $\mathcal{C}^{\infty}$ on $(0,s)\times\mathcal{O}$, no matter the initial condition $u_{0}$.
B.2. Background on Diffusion Markov Triples
We first introduce Diffusion Markov Triples.
Definition 18 (Diffusion Markov Triple).
A Diffusion Markov Triple $(E,\mu,\Gamma)$ is composed of a measure space $(E,\mathcal{F},\mu)$, a class $\mathcal{A}_{0}$ of bounded measurable functionals on $E$, and a symmetric bilinear operator (the carré du champ operator) $\Gamma:\mathcal{A}_{0}\times\mathcal{A}_{0}\rightarrow\mathcal{A}_{0}$, satisfying properties D1-D9 in Bakry et al. (2014, Section 3.4.1). We report them here to keep the paper self-contained.
  • (DMT1)
    Measure space $(E,\mathcal{F},\mu)$ is a good measure space, that is, there is a countable family of sets which generates $\mathcal{F}$ (up to sets of $\mu$-measure 0), and both the decomposition and bi-measure theorems apply. Measure $\mu$ is $\sigma$-finite and, when finite, assumed to be a probability.
  • (DMT2)
    $\mathcal{A}_{0}$ is a vector space of bounded measurable functionals on $E$, which is dense in every $\mathbb{L}^{p}(\mu)$, $1\leq p<\infty$, stable under products (that is, $\mathcal{A}_{0}$ is an algebra) and stable under the action of smooth ($\mathcal{C}^{\infty}$) functions $\Psi:\mathbb{R}^{k}\rightarrow\mathbb{R}$ vanishing at 0, i.e. such that $\Psi(0)=0$.
  • (DMT3)
    The carré du champ operator $\Gamma$ is positive, that is, $\Gamma(f)\equiv\Gamma(f,f)\geq 0$, for all $f\in\mathcal{A}_{0}$. [^30] The carré du champ operator $\Gamma$ also satisfies Bakry et al. (2014, Equation (3.1.1)), that is, $\int_{E}\Gamma(g,f^{2})\text{d}\mu+2\int_{E}g\Gamma(f)\text{d}\mu=2\int_{E}\Gamma(fg,f)\text{d}\mu$. It satisfies the following diffusion hypothesis: for every smooth ($\mathcal{C}^{\infty}$) function $\Psi$ vanishing at 0, and every $f_{1},\ldots,f_{k},g\in\mathcal{A}_{0}$,
    $\Gamma(\Psi(f_{1},\ldots,f_{k}),g)=\sum_{i=1}^{k}\partial_{i}\Psi(f_{1},\ldots,f_{k})\Gamma(f_{i},g).$ (21)
  • (DMT4)
    For every $f\in\mathcal{A}_{0}$, there exists a finite constant $C(f)$ such that for every $g\in\mathcal{A}_{0}$,
    $\left|\int_{E}\Gamma(f,g)\text{d}\mu\right|\leq C(f)\|g\|_{2}.$
    The Dirichlet form $\mathcal{E}$ is defined for every $(f,g)\in\mathcal{A}_{0}\times\mathcal{A}_{0}$ by
    $\mathcal{E}(f,g)=\int_{E}\Gamma(f,g)\text{d}\mu,$
    and we write $\mathcal{E}(f,f)\equiv\mathcal{E}(f)$ for notational convenience. The domain $\mathcal{D}(\mathcal{E})$ of the Dirichlet form $\mathcal{E}$ is the completion of $\mathcal{A}_{0}$ with respect to the norm $\|f\|_{\mathcal{E}}=[\|f\|_{2}^{2}+\mathcal{E}(f)]^{1/2}$. The Dirichlet form $\mathcal{E}$ is extended to $\mathcal{D}(\mathcal{E})$ by continuity together with the carré du champ operator $\Gamma$.
  • (DMT5)
    $L$ is a linear operator on $\mathcal{A}_{0}$ defined by and satisfying the integration by parts formula
    $\int_{E}gLf\text{d}\mu=-\int_{E}\Gamma(f,g)\text{d}\mu,$
    for all $f,g\in\mathcal{A}_{0}$. A consequence of condition (DMT3) is that
    $\begin{split}L\left(\Psi(f_{1},\ldots,f_{k})\right)&=\sum_{i=1}^{k}\partial_{i}\Psi(f_{1},\ldots,f_{k})Lf_{i}\\ &+\sum_{i,j=1}^{k}\partial^{2}_{ij}\Psi(f_{1},\ldots,f_{k})\Gamma(f_{i},f_{j}),\end{split}$ (22)
    for a smooth ($\mathcal{C}^{\infty}$) functions $\Psi:\mathbb{R}^{k}\rightarrow\mathbb{R}$ such that $\Psi(0)=0$, and $f_{1},\ldots,f_{k}\in\mathcal{A}_{0}$. Hence $L$ too satisfies the diffusion property as we introduced it in Definition 15.
  • (DMT6)
    For the operator L defined in (DMT5), $L(\mathcal{A}_{0})\subset\mathcal{A}_{0}$, i.e. $Lf\in\mathcal{A}_{0}$, for all $f\in\mathcal{A}_{0}$.
  • (DMT7)
    The domain $\mathcal{D}(L)$ of the operator $L$ is defined as the set of $f\in\mathcal{D}(\mathcal{E})$ for which there exists a finite constant $C(f)$ such that, for any $g\in\mathcal{D}(\mathcal{E})$,
    $\left|\mathcal{E}(f,g)\right|\leq C(f)\|g\|_{2}.$
    On $\mathcal{D}(L)$, $L$ is extended via the integration by parts formula for every $g\in\mathcal{D}(\mathcal{E})$.
  • (DMT8)
    For every $f\in\mathcal{A}_{0}$, $\int_{E}Lf\text{d}\mu=0$.
  • (DMT9)
    The semigroup $\mathbf{P}=(P_{t})_{t\geq 0}$ is the symmetric semigroup with infinitesimal generator $L$ defined on its domain $\mathcal{D}(L)$. In general, it is sub-Markov, that is, $P_{t}(\mathbbm{1})\leq\mathbbm{1}$, for all $t\geq 0$, where $\mathbbm{1}$ is the constant function equal to $1$. For it to be Markov, $P_{t}(\mathbbm{1})=\mathbbm{1}$ must hold for all $t\geq 0$. Moreover, $\mathcal{A}_{0}$ is not necessarily dense in the domain $\mathcal{D}(L)$ with respect to the domain norm $\|f\|_{\mathcal{D}(L)}=[\|f\|_{2}^{2}+\|Lf\|_{2}^{2}]^{1/2}$. By (DMT8), we have that $\mu$ is invariant for $\mathbf{P}$. In addition, from the symmetry of $\Gamma$ we have that $\mu$ is reversible for $\mathbf{P}$.
Given Definition 18, we see how $L$ is a second order differential operator, $\Gamma$ is a first order operator in each of its arguments, and $\Gamma_{2}$ (see Definition 16) is a second order differential in each of its arguments.
We now introduce the concepts of adjoint operator, self-adjointness and essential self-adjointness (ESA) (Bakry et al., 2014, Section 3.4.2). The latter allow us in Theorem 10, to relax the requirement in Theorem 8 that $\underline{L}$ and $\overline{L}$ are elliptic operators.
Definition 19 (Adjoint Operator).
Let $(E,\mu,\Gamma)$ be a diffusion Markov triple, and consider the operator $L$ associated with it. The domain $\mathcal{D}(L^{*})$ is the set of functions $f\in\mathbb{L}^{2}(\mu)$ such that there exists a finite constant $C(f)$ for which, for all $g\in\mathcal{A}_{0}$,
$\left|\int_{E}fLg\text{d}\mu\right|\leq C(f)\|g\|_{2}.$
On this domain, the adjoint operator $L^{*}$ is defined by integration by parts: for any $g\in\mathcal{A}_{0}$,
$L^{*}(f)=\int_{E}fLg\text{d}\mu=\int_{E}gL^{*}f\text{d}\mu.$
By symmetry of $L$ – as we have seen in (16) – it holds that (in general) $\mathcal{D}(L)\subset\mathcal{D}(L^{*})$, and $L=L^{*}$ on $\mathcal{D}(L)$. That is, $L$ is self-adjoint on $\mathcal{D}(L)$. Being self-adjoint on $\mathcal{D}(L)$, though, does not mean that $L=L^{*}$. The self-adjointness property of $L$ refers to the dual operator constructed in the same way as $L^{*}$, but where $\mathcal{D}(L)$ replaces $\mathcal{A}_{0}$. In other words, as we have seen in Definition 18.(DMT9), $\mathcal{A}_{0}$ is not necessarily dense in $\mathcal{D}(L)$ with respect to the domain norm $\|\cdot\|_{\mathcal{D}(L)}$. This happens if and only if $\mathcal{D}(L)=\mathcal{D}(L^{*})$. In this case, the extension of $L$ from $\mathcal{A}_{0}$ to a larger domain as a self-adjoint operator is unique.
Definition 20 (Essential Self-Adjointness – ESA).
The operator $L$ is said to be essentially self-adjoint if $\mathcal{D}(L)=\mathcal{D}(L^{*})$ (recall that $\mathcal{D}(L^{*})$ is defined with respect to $\mathcal{A}_{0}$). Equivalently, $\mathcal{A}_{0}$ is dense in $\mathcal{D}(L)$ in the topology induced by $\|\cdot\|_{\mathcal{D}(L)}$.
As we can see, the ESA property depends on the choice of $\mathcal{A}_{0}$. In addition, ESA holds for all elliptic operators on complete Riemannian manifolds where $\mathcal{A}_{0}$ is the class of smooth compactly supported functions. That is, it is always satisfied when $E$ is a Euclidean space or a complete Riemannian manifold. The other implication, though, need not hold – that is, there are ESA operators that are not elliptic as in B.1.1. As we shall see, ESA alone is enough to guarantee ergodicity à la Corollary 8.1.
The ESA property does not automatically hold for a diffusion Markov triple $(E,\mu,\Gamma)$. To derive sufficient conditions for it, we need to introduce the concept of an extension of $\mathcal{A}_{0}$ (Bakry et al., 2014, Section 3.3.1).
Definition 21 (Extended Algebra $\mathcal{A}$).
Let $\mathcal{A}_{0}$ be the algebra associated with a diffusion Markov triple $(E,\mu,\Gamma)$. $\mathcal{A}$ is an algebra of measurable functions on $E$ containing $\mathcal{A}_{0}$, containing the constant functions, and satisfying the following requirements.
  • (i)
    For all $f\in\mathcal{A}$ and all $h\in\mathcal{A}_{0}$, $hf\in\mathcal{A}_{0}$.
  • (ii)
    For any $f\in\mathcal{A}$, if $\int_{E}hf\text{d}\mu\geq 0$ for all positive $h\in\mathcal{A}_{0}$, then $f\geq 0$.
  • (iii)
    $\mathcal{A}$ is stable under composition with smooth ($\mathcal{C}^{\infty}$) functions $\Psi:\mathbb{R}^{k}\rightarrow\mathbb{R}$.
  • (iv)
    The operator $L:\mathcal{A}\rightarrow\mathcal{A}$ is an extension of the infinitesimal generator $L$ on $\mathcal{A}_{0}$. The carré du champ operator $\Gamma$ is also defined on $\mathcal{A}\times\mathcal{A}$ as in Definition 14, with respect to the extension of $L$.
  • (v)
    For every $f\in\mathcal{A}$, $\Gamma(f)\geq 0$.
  • (vi)
    The operators $\Gamma$ and $L$ satisfy (21) and (22), respectively, for any smooth ($\mathcal{C}^{\infty}$) functions $\Psi:\mathbb{R}^{k}\rightarrow\mathbb{R}$.
  • (vii)
    For every $f\in\mathcal{A}$ and every $g\in\mathcal{A}_{0}$, the following holds
    $\mathcal{E}(f,g)=\int_{E}\Gamma(f,g)\text{d}\mu=-\int_{E}gLf\text{d}\mu=-\int_{E}fLg\text{d}\mu.$
  • (viii)
    For every $f\in\mathcal{A}_{0}$ and every $t\geq 0$, $P_{t}f\in\mathcal{A}$.
When $E$ is a Euclidean space or a complete Riemannian manifold, we have that $\mathcal{A}_{0}$ is the class of smooth ($\mathcal{C}^{\infty}$) compactly supported functionals on $E$, while $\mathcal{A}$ is the class of smooth ($\mathcal{C}^{\infty}$) functionals. We now present two concepts, connexity and weak hypo-ellipticity, that allow to “import” some of the geometric structure of the Riemannian manifolds considered in B.1.1 to the more general diffusion Markov triple setting (Bakry et al., 2014, Section 3.3.3). In addition, as we can see in the proof of Theorem 10, in the context of imprecise diffusion Markov tuples they are sufficient for the ESA property to hold.
Definition 22 (Connexity and Weak Hypo-Ellipticity).
Let $(E,\mu,\Gamma)$ be a diffusion Markov triple, and let $\mathcal{A}$ be the extension of the algebra $\mathcal{A}_{0}$ associated with $(E,\mu,\Gamma)$. Then, $(E,\mu,\Gamma)$ is said to be connected if $f\in\mathcal{A}$ and $\Gamma(f)=0$ imply that $f$ is constant. This is a local property for functionals in $\mathcal{A}$. $(E,\mu,\Gamma)$ is weakly hypo-elliptic if, for every $\lambda\in\mathbb{R}$, any $f\in\mathcal{D}(L^{*})$ satisfying $L^{*}f=\lambda f$ belongs to $\mathcal{A}$.
Notice also that in this more general setting, the Bakry-Émery curvature condition is the same as that in Definition 17, where the sufficiently rich class corresponds to the extended algebra $\mathcal{A}$.